# Complementary Search of Fermionic Absorption Operators at Hadron Collider and Direct Detection Experiments

### v2

@@ Abstract
Instead of the energy recoil signal at direct detection experiments, dark fermion appears always as missing energy at high energy colliders. For a fermionc dark sector particle coupled with quarks and neutrino via absorption operators, its production is always accompanied by an invisible neutrino. We study in details the mono-X (photon, jet, and $Z$) productions at the Large Hadron Collider (LHC). To make easy comparison between the collider and direct detection experiments, we start from the quark-level absorption operators for the first time. In other words, we study the model-independent constraints on a generic fermionic absorption dark fermion. In addition, the interplay and comparison with the possible detection at the neutrino experiment especially Borexino is also briefly discussed. We find that for both spin-dependent and spin-independent absorption of the dark fermion, the experiments with light nuclear target can provide the strongest constraints.
@@ 1 Introduction
With high center-of-mass energy, collider search has an additional superiority: it can even probe the heavy DS particles that are usually contained in a UV model [16, 17, 18, 20, 19], as long as its direct production is kinematically allowed. The collider search can also provide a controllable environment to distinguish the leptophilic and hadrophilic nature of the DS couplings with SM particles. While hadron colliders mainly probe quark couplings to DS particles [57, 58, 59, 60, 61], lepton colliders are more sensitive to the leptonic ones [62, 63, 64, 65, 66, 67, 68, 69, 70, 15]. Furthermore, the isospin violation and lepton universality of the DS-quark and DS-lepton couplings can also be studied at hadron and lepton colliders, respectively. Therefore, collider search of the DS particles is not just complementary but also unique for DS particles (or WIMP) [52, 53, 54, 71, 72, 73, 74, 15, 61].
@@ 2 Fermionic Dark Particle Absorption Operators
In this paper, we study the interactions that can induce the absorption of fermionic DS particle by nuclei [11, 12, 96, 97, 98, 99, 100, 61]. Instead of transferring the kinetic energy of the fermionic DS particle into the nuclei recoil energy by elastic scattering, in the absorption process a fermionic DS particle is converted to a neutrino. Both its mass and kinetic energy are transferred into the nuclei recoil energy. Such processes can be effectively initiated by four-fermion contact interactions [11, 12, 100, 61] which is very similar to the absorption process at electron target [13, 14, 99, 101, 102, 15, 103] .
In the effective field theory (EFT) framework, the interaction operators between the DS particles and the SM particles are usually constructed according to the SM $SU_{C}(3)\times SU_{L}(2)\times U_{Y}(1)$ gauge symmetries. Assuming that the isospin symmetry is preserved and denoting the first generation up ($u$) and down ($d$) quarks as an isospin doublet $q=(u~{}d)^{T}$, the relevant leading order dim-6 operators connecting the quark and DS paticle-neutrino current can be written as,
@@
as well as their hermitian conjugates. The quark currents $\bar{q}\varGamma q$ with $\varGamma\equiv 1$, $i\gamma_{5}$, $\gamma_{\mu}$, $\gamma_{\mu}\gamma_{5}$ and $\sigma_{\mu\nu}$ should be understood as $\bar{q}\varGamma q=\bar{u}\varGamma u+\bar{d}\varGamma d$ from the isospin symmetry assumption. The above parameterization includes all the five independent Lorentz structures of the quark bilinear and is complete in the sense that any other dim-6 operator can be written as a linear combination of the above 5 operators by employing $\gamma$-algebra and Fierz transformations [104, 105, 106]. The neutrino is assumed to be left-handed while the dark matter $\chi$ is a Dirac fermion with both left- and right-handed components. The effective Lagrangian takes the form as,
[…]
@@
The four-fermion contact interactions in Eq. (2.1) can lead to the associated production of the DS particle and neutrino at hadron collider. The above EFT description of collider searches can work well as long as the cut-off scale $\varLambda_{i}$ [49] or the mediator mass [107, 108, 109] is much larger than the collision energy. On the other hand, because of the running effect, the EFT operators defined at the TeV scale can become a mixture at low energy [110, 111, 112, 113, 114, 52, 53, 54]. However, the mixing effect of the effective operators can be numerically relevant only when the DS particle couples with the top quark [53]. At hadron colliders, the tree-level direct production of the DS particle is sensitive to only the light quarks. Here we assume that the effective operators in Eq. (2.1) is only valid for light quarks and hence the mixing effect can be safely neglected. The interaction between the DS particle and top quark can be searched for via the associated production of top quark(s) and DS particle at hadron collider while its influences on the running effect can be investigated consistently. We will study this part elsewhere.
@@ 3.1 Mono-Photon
The panels (a) and (b) of Fig. 2 show the photon normalized polar angle $\theta_{\gamma}$ and transverse momentum $p_{T,\gamma}$ distributions in the laboratory frame at LHC-13 ($\sqrt{s}=13\,{\rm TeV}$ and a total luminosity $\mathcal{L}=139\,{\rm fb}^{-1}$). The signal properties (colorful non-solid curves) are illustrated with $m_{\chi}=0\,{\rm GeV}$ and $\varLambda_{i}=1\,{\rm TeV}$. As expected, the signal events are dominant at the forward and backward regions which is similar as the irreducible background distribution (black-solid curve). We can also see in Fig. 2 (b) that both the signal and background events mostly distribute in the soft $p_{T,\gamma}$ region. However, the signal distributions are larger than its background counterpart for $p_{T,\gamma}\gtrsim 200\,{\rm GeV}$. Hence $p_{T,\gamma}=200\,{\rm GeV}$ is a good cut to enhance the signal significance. Fig. 2 (c) shows the dependence of the signal and background total parton-level cross sections on the center-of-mass energy $\sqrt{\hat{s}}$. While the background decreases quickly with increasing $\sqrt{\hat{s}}$, the signals grow rapidly and exceed the background around $\sqrt{\hat{s}}=(700\sim 800)\,{\rm GeV}$ which is less than the energy cutoff $\varLambda_{i}=1\,{\rm TeV}$.
@@
The mono-$\gamma$ process has been extensively studied at LHC for various DS particle models containing a mediator [115]. It is expected that the same data can give exclusion limits on our model at the same level, i.e., $\varLambda_{i}\sim m_{\rm med}$. Our simulations are conducted using the toolboxes MadGraph [116, 117] and FeynRules [118]. The signal total cross sections are estimated according to the event selections given in the ATLAS search [115]. A photon with a transverse energy above 130 GeV was required at the matrix-element level. A strong kinematical cut ${E\!\!\!/}_{T}\gtrsim 200\,{\rm GeV}$ is used to select the signal events. In this region, the parton shower effect is negligible. So our simulation is done at the generator level. The total detection efficiency is accounted for by an overall normalization factor which is estimated by validating the irreducible background process $pp\to\gamma Z(\nu\nu)$. Consequently, the detector-level predictions are obtained by multiplying this normalization factor with the generator-level results.
@@ 3.3 Mono-$Z$
However, it is not straightforward to study the experimental constraints on the model parameters, because the $Z$ reconstruction is necessary in practical measurement. For its leptonic decay modes, since tracks can be measured precisely, the mono-$Z$ events can be efficiently selected by putting a cut on the lepton pair invariant mass. However, it is not true for the hadronic decay modes. On one hand, the jet momentum uncertainty is much larger than its lepton counterpart. On the other hand, both the electroweak (for instance $pp\to W^{\pm}(jj)+{E\!\!\!/}_{T}$) and the pure QCD channels $pp\to jj+{E\!\!\!/}_{T}$ can contribute as background. Actually, the pure QCD contribution completely dominates the total background [120]. So we study the leptonic and hadronic decay modes separately in Sec. 3.3.1 and Sec. 3.3.2.
@@ 3.3.1 Leptonic decay modes
Let us first study the mono-$Z$ production followed by leptonic decays $Z\to\ell\ell$ ($\ell=e$ and $\mu$) at the center-of-mass energy $\sqrt{s}=13\,{\rm TeV}$. The left and right panels of Fig. 9 show the normalized polar angle and transverse momentum distributions of the reconstructed $Z$-boson from the two leptons. One can clearly see that for both the signal (colorful non-solid curves) and the irreducible background (black-solid curve), the distinctive kinematic properties (that has been discussed for the on-shell $Z$ production) of the mono-$Z$ events can be readily reconstructed.
@@ 3.3.2 Hadronic decay modes
The mono-$Z$ event reconstruction for the hadronic decay mode is much more complex than the leptonic one.
[…]
Firstly, because of the relatively small mass difference between the $W$ and $Z$ bosons as well as the large jet momentum uncertainty, the reconstructed $Z$ bosons are inevitably contaminated by the hadronic $W$ decay products. Fig. 12 shows the normalized polar angle $\cos\theta_{V(jj)}$ and transverse momentum $p_{T,V(jj)}$ distributions of the reconstructed vector boson from $pp\to V(jj)+{E\!\!\!/}_{T}$ for both $V=Z$ and $W$. Comparing with the purely mono-$Z$ contribution in Fig. 8 (a), the background is less towards the forward and backward regions. This is due to the large jet transverse energy cut, as shown in Fig. 12 (b).
[…]
The ATLAS collaboration has searched for dark matter in the mono-$Z/W$ production with hadronically decaying $Z/W$-boson [120] in three categorizations: the high purity (HP) and low purity (LP) regions for merged topology (MT), and the resolved topology (RT). In each region, 3 configurations are defined according to $b$-tagging. Since the background in the RT region is about 1 order larger, our following calculations consider only the LP and HP regions. Furthermore, the contributions with non-zero $b$-tagging number are sub-leading and hence can be ignored.
@@
Fig. 14 compares our simulation (red triangle) and the ATLAS result (black-solid line) for the missing transverse energy distribution of the pure QCD di-jet production channel $pp\to Z(\nu\nu)+jj$ at the LHC-13 with a total luminosity $\mathcal{L}=36.1\,{\rm fb}^{-1}$. The total event number of our result has been normalized to the experimental data. One can see that the detector effects can be properly described by an overall normalization factor. The same normalization factor is then multiplied with the signal cross sections to calculate the corresponding exclusion limits.
@@ 3.4 Projected Sensitivities at Future Upgrades
${E\!\!\!/}_{T}\geqslant 400\,\,{\rm GeV}$
$pp\to Z(\ell^{+}\ell^{-}){E\!\!\!/}_{T}$
[…]
@@
Fig. 16 and Fig. 17 show the expected 95% C.L. exclusion limits using the mono-$\gamma$, mono-$j$, mono-$Z_{\ell\ell}$, and mono-$V_{jj}$ (combined results of the HP and LP regions) events at the LHC-14 (left) and LHC-25 (right), respectively.
@@
One can clearly see the constraint improvements, although the enhancements vary from process to process. However, for all the processes, there is roughly a factor of 1.5 and 3 enhancement on the lower bounds of the energy scale $\Lambda_{i}$ (for $m_{\chi}=0$) at the LHC-14 and LHC-25 configurations, respectively.

### v3

@@ Abstract
Instead of the energy recoil signal at direct detection experiments, dark fermion appears as missing energy at hadron colliders. For a fermionc dark sector particle that coupled with quarks and neutrino via absorption operators, its production at collider is accompanied by an invisible neutrino. We study in details the mono-$X$ (photon, jet, and $Z$) productions at the Large Hadron Collider (LHC). We start from the quark-level absorption operators to make easy comparison between the collider and direct detection experiments. In other words, we study the model-independent constraints on a dark fermion with absorption operator. In addition, the interplay and comparison with the possible detection at the neutrino experiments, especially Borexino, is also briefly discussed. We find that light nuclear target can provide the stronger constraints on both spin-dependent and spin-independent absorption operators.
@@ 1 Introduction
With high center-of-mass energy, collider search has an additional superiority: it can even probe the heavy DS particles that are usually contained in a UV model [16, 17, 18, 20, 19], as long as its direct production is kinematically allowed. The collider search can also provide a controllable environment to distinguish the leptophilic and hadrophilic nature of the DS couplings with SM particles. While hadron colliders mainly probe quark and gluon couplings to DS particles [57, 58, 59, 60, 61], lepton colliders are more sensitive to the leptonic ones [63, 64, 65, 66, 67, 68, 69, 70, 71, 15]. Furthermore, the isospin violation and lepton universality of the DS-quark and DS-lepton couplings can also be studied at hadron and lepton colliders, respectively. Therefore, collider search of the DS particles is not just complementary but also unique for DS particles (or WIMP) [52, 53, 54, 72, 73, 74, 75, 15, 61].
@@ 2 Fermionic Dark Particle Absorption Operators
In this paper, we study the interactions that can induce the absorption of a fermionic DS particle by nuclei [11, 12, 97, 98, 99, 100, 101, 61]. Instead of transferring the kinetic energy of a fermionic DS particle into the nuclei recoil energy by elastic scattering, the fermionic DS particle is converted to a neutrino in the absorption process. Both its mass and kinetic energy are transferred into the nuclei recoil and neutrino energies. Such process can be effectively initiated by four-fermion contact interactions [11, 12, 101, 61] which is very similar to the absorption process on an electron target [13, 14, 100, 102, 103, 15, 104] .
In the effective field theory (EFT) framework, the interaction operators between the DS and SM particles are usually constructed according to the SM $SU_{C}(3)\times SU_{L}(2)\times U_{Y}(1)$ gauge symmetries. Assuming that the isospin symmetry is preserved and denoting the first generation up ($u$) and down ($d$) quarks as an isospin doublet $q\equiv(u~d)^{T}$, the relevant leading order dim-6 operators connecting the quark and DS paticle-neutrino current can be written as,
@@
as well as their hermitian conjugates. Although we put the left-handed neutrinos here, it is equivalent to replace by its right-handed counterpart. The quark currents $\bar{q}\varGamma q$ with $\varGamma\equiv 1$, $i\gamma_{5}$, $\gamma_{\mu}$, $\gamma_{\mu}\gamma_{5}$ and $\sigma_{\mu\nu}$ should be understood as $\bar{q}\varGamma q=\bar{u}\varGamma u+\bar{d}\varGamma d$ from the isospin symmetry assumption. The above parameterization includes all the five independent Lorentz structures of the quark bilinear and is complete in the sense that any other dim-6 operator can be written as a linear combination of the above 5 operators by employing the $\gamma$ algebra and Fierz transformations [105, 106, 107]. The neutrino is assumed to be left-handed while the dark matter $\chi$ is a Dirac fermion with both left- and right-handed components. The effective Lagrangian takes the form as,
$\mathcal{L}_{\text{eff }}=\sum_{i}\frac{1}{\varLambda_{i}^{2}}\mathcal{O}_{i}+\text{h.c.},$ (2.2)
@@
The four-fermion contact interactions in Eq. (2.1) can lead to the associated production of DS particle and neutrino at hadron collider. The above EFT description of collider searches can work well as long as the cut-off scale $\varLambda_{i}$ [49] or the mediator mass [108, 109, 110] is much larger than the collision energy. More concretely, suppose that couplings of the DS paticle $\chi$ and the quarks to a dark vector $V^{\prime}$ of a dark $U(1)$ gauge group are $g_{\chi}$ and $g_{q}$, respectively, and the mixing strength of $\chi$ and neutrino as induced by a dark scalar is $\varTheta$, then one has $m_{V^{\prime}}=\varTheta g_{\chi}g_{q}\varLambda$ for the heavy dark mediator $V^{\prime}$ [12]. Requiring $m_{V^{\prime}}\gtrsim\varLambda$ implies $\varTheta g_{\chi}g_{q}\gtrsim 1$. For a conservative estimation, we simply require $g_{\chi}\sim g_{q}\sim 4\pi$ for validation of perturbativity. Then, the mixing strength parameter can be much smaller, $\varTheta\gtrsim 6.3\times 10^{-3}$. Since such a weak mixing effect have not been excluded experimentally, our effective operators descriptions are still valid for a conservative estimation of the exclusion limits. Models with lighter mediator ($\varLambda$) can get more stringent constraints.
Because of the running effect, the EFT operators defined at the TeV scale can become a mixture at low energy [111, 112, 113, 114, 115, 52, 53, 54]. However, the mixing effect of the effective operators can be numerically relevant only when the DS particle couples with the top quark [53]. At hadron colliders, the tree-level direct production of the DS particle is sensitive to only the light quarks. Here we assume that the effective operators in Eq. (2.1) is only valid for light quarks and hence the mixing effect can be safely neglected. The interaction between the DS particle and top quark can be searched for via the associated production of top quark(s) and DS particle at hadron collider while its influences on the running effect can be investigated consistently. We will study this part elsewhere.
We would like to point out again that the DS particles need not to be the DM that survives until now in the Universe. Any DS fermion can have absorption operator with neutrino. Although such operator can lead to decay of the DS fermion, its stability is not a requirement if they are not DM. For a dark fermion mass with $m_{\chi}\lesssim 5\ \text{GeV}$, a variety of thermal and non-thermal production mechanisms for such light dark fermions can accommodate the observed dark matter relic abundance [11, 12]. In this paper, rather than elaborating on these production mechanisms, we adopt a model-independent approach to study signal properties at hadron colliders and direct detection experiments. A collider search can probe not just DM but the DS particles in general as detailed below in Sec. 3. The complementary search at direct detection experiments for a DM particle with small enough mass to guarantee its stability can be found in Sec. 4.
@@ 3.1 Mono-Photon
The panels (a) and (b) of Fig. 2 show the photon normalized polar angle $\theta_{\gamma}$ and transverse momentum $p_{T,\gamma}$ distributions in the laboratory frame at LHC-13 ($\sqrt{s}=13\,{\rm TeV}$ and a total luminosity $\mathcal{L}=139\,{\rm fb}^{-1}$). Although the collider search can scan a much wider mass range, we choose $m_{\chi}=0\,{\rm GeV}$ and $\varLambda_{i}=1\,{\rm TeV}$ to illustrate the signal properties (colorful non-solid curves). For a small enough DM mass, $m_{\chi}\ll\sqrt{\hat{s}}$ where $\sqrt{\hat{s}}$ is the effective center-of-mass energy in the primary scattering process, the event cross section and topology would become essentially insensitive to the DM mass. Then, the illustrated results with vanishing mass can apply more generally. As expected, the signal events are dominant at the forward and backward regions which is similar as the irreducible background distribution (black-solid curve). We can also see in Fig. 2 (b) that both the signal and background events mostly distribute in the soft $p_{T,\gamma}$ region. However, the signal distributions are larger than its background counterpart for $p_{T,\gamma}\gtrsim 200\,{\rm GeV}$. Hence $p_{T,\gamma}=200\,{\rm GeV}$ is a good cut to enhance the signal significance. Fig. 2 (c) shows the dependence of the signal and background total parton-level cross sections on the center-of-mass energy $\sqrt{\hat{s}}$. While the background decreases quickly with increasing $\sqrt{\hat{s}}$, the signals grow rapidly and exceed the background around $\sqrt{\hat{s}}=(700\sim 800)\,{\rm GeV}$ which is less than the energy cutoff $\varLambda_{i}=1\,{\rm TeV}$.
The mono-$\gamma$ process has been extensively studied at LHC for various DS particle models containing a mediator [116]. It is expected that the same data can give exclusion limits on our model at the same level, i.e., $\varLambda_{i}\sim m_{\rm med}$. Our simulations are conducted using the toolboxes MadGraph [117, 118] and FeynRules [119]. The signal total cross sections are estimated according to the event selections given in the ATLAS search [116]. A photon with a transverse energy above 130 GeV was required at the matrix-element level. A strong kinematical cut ${E\!\!\!/\penalty 50}_{T}\gtrsim 200\,{\rm GeV}$ is used to select the signal events. In this region, the parton shower effect is negligible. So our simulation is done at the generator level. The total detection efficiency is accounted for by an overall normalization factor which is estimated by validating the irreducible background process $pp\to\gamma Z(\nu\nu)$. Consequently, the detector-level predictions are obtained by multiplying this normalization factor with the generator-level results.
@@ 3.3 Mono-$Z$
However, it is not straightforward to study the experimental constraints on the model parameters, because the $Z$ reconstruction is necessary in practical measurement. For its leptonic decay modes, since tracks can be measured precisely, the mono-$Z$ events can be efficiently selected by putting a cut on the lepton pair invariant mass. However, it is not true for the hadronic decay modes. On one hand, the jet momentum uncertainty is much larger than its lepton counterpart. On the other hand, both the electroweak (for instance $pp\to W^{\pm}(jj)+{E\!\!\!/\penalty 50}_{T}$) and the pure QCD channels $pp\to jj+{E\!\!\!/\penalty 50}_{T}$ can contribute as background. Actually, the pure QCD contribution completely dominates the total background [121]. So we study the leptonic and hadronic decay modes separately in Sec. 3.3.1 and Sec. 3.3.2.
@@ 3.3.1 Leptonic decay modes
Let us first study the mono-$Z$ production followed by leptonic decays $Z\to\ell\ell$ ($\ell=e$ and $\mu$) at the center-of-mass energy $\sqrt{s}=13\,{\rm TeV}$.
@@
The left and right panels of Fig. 9 show the normalized polar angle and transverse momentum distributions of the reconstructed $Z$-boson from the two leptons. One can clearly see that for both the signal (colorful non-solid curves) and the irreducible background (black-solid curve), the distinctive kinematic properties (that has been discussed for the on-shell $Z$ production) of the mono-$Z$ events can be readily reconstructed.
@@ 3.3.2 Hadronic decay modes
The mono-$Z$ event reconstruction for the hadronic decay mode is much more complex than the leptonic one. Firstly, because of the relatively small mass difference between the $W$ and $Z$ bosons as well as the large jet momentum uncertainty, the reconstructed $Z$ bosons are inevitably contaminated by the hadronic $W$ decay products. Fig. 12 shows the normalized polar angle $\cos\theta_{V(jj)}$ and transverse momentum $p_{T,V(jj)}$ distributions of the reconstructed vector boson from $pp\to V(jj)+{E\!\!\!/\penalty 50}_{T}$ for both $V=Z$ and $W$. Comparing with the purely mono-$Z$ contribution in Fig. 8 (a), the background is less towards the forward and backward regions. This is due to the large jet transverse energy cut, as shown in Fig. 12 (b).
Fig. 12: The normalized polar angle (left panel) and transverse momentum (right panel) distributions of the reconstructed vector boson from two jets with invariant mass $m_{jj}\in[65,\,105]\,{\rm GeV}$ for two channels $pp\to V(jj)+{E\!\!\!/\penalty 50}_{T}$ ($V=Z$ and $W$) at center-of-mass energy $\sqrt{s}=13\,{\rm TeV}$. In both panels, the signal events with parameters $m_{\chi}=0\,{\rm GeV}$ and $\varLambda_{i}=1\,{\rm TeV}$ are illustrated by colorful non-solid curves and the background is shown as black-solid curve.
Fig. 13: The normalized polar angle (left) and transverse momentum (right) distributions in the invariant mass window $m_{\rm jets}\in[65,\,105]\,{\rm GeV}$ for the channels $pp\to{E\!\!\!/\penalty 50}_{T}+{\rm jets}$ with QCD jets, respectively. For all the panels, the signal events (colorful non-solid curves) are illustrated with parameters $m_{\chi}=0\,{\rm GeV}$ and $\varLambda_{i}=1\,{\rm TeV}$.
Secondly, the mono-$Z$ events are even more heavily contaminated by the associated production of missing energy with pure QCD jets. The dominant contribution comes from the associated production of 2-jets with an invisibly decaying $Z$ boson, i.e., the channels $pp\to Z(\nu\nu)+jj$. In this case, the 2 jets with invariant mass in the window $m_{\rm jets}\in[65,\,105]\,{\rm GeV}$ can be miss-identified as a hadronically decaying $Z$ boson. Actually, it consists majority of the total background [121]. Fig. 13 shows the polar angle and transverse momentum distributions of the summed jet ($jj$) for both the signal processes $pp\to\chi\nu+2j$ and the dominant background $pp\to Z(\nu\nu)+2j$. Comparing with Fig. 12, both signal and background are pushed towards the forward and backward regions. Consequently, the transverse momentum of the fake $Z$-boson from QCD jets are softer.
The ATLAS collaboration has searched for dark matter in the mono-$Z/W$ production with hadronically decaying $Z/W$-boson [121] in three categorizations: the high purity (HP) and low purity (LP) regions for merged topology (MT), and the resolved topology (RT). In each region, 3 configurations are defined according to $b$-tagging. Since the background in the RT region is about 1 order larger, our following calculations consider only the LP and HP regions. Furthermore, the contributions with non-zero $b$-tagging number are sub-leading and hence can be ignored. Fig. 14 compares our simulation (red triangle) and the ATLAS result (black-solid line) for the missing transverse energy distribution of the pure QCD di-jet production channel $pp\to Z(\nu\nu)+jj$ at the LHC-13 with a total luminosity $\mathcal{L}=36.1\,{\rm fb}^{-1}$. The total event number of our result has been normalized to the experimental data. One can see that the detector effects can be properly described by an overall normalization factor. The same normalization factor is then multiplied with the signal cross sections to calculate the corresponding exclusion limits.
@@ 3.4 Projected Sensitivities at Future Upgrades
${E\!\!\!/\penalty 50}_{T}\geqslant 400\,\,{\rm GeV}$
$pp\to Z(\ell^{+}\ell^{-}){E\!\!\!/\penalty 50}_{T}$
${E\!\!\!/\penalty 50}_{T}\geqslant\,85\,\,{\rm GeV}$
@@
Fig. 16 and Fig. 17 show the expected 95% C.L. exclusion limits using the $\gamma, mono$-$j$, mono-$Z_{\ell\ell}$, and mono-$V_{jj}$ (combined results of the HP and LP regions) events at the LHC-14 (left) and LHC-25 (right), respectively. One can clearly see the constraint improvements, although the enhancements vary from process to process. However, for all the processes, there is roughly a factor of 1.5 and 3 enhancement on the lower bounds of the energy scale $\Lambda_{i}$ (for $m_{\chi}=0$) at the LHC-14 and LHC-25 configurations, respectively.
3.5 Possible Enhancements
The sensitivities obtained above are very conservative in two aspects. First, in the mixing picture provided by the UV completion [12], not just those absorption operators in Eq. (2.1) with one dark fermion $\chi$ and one neutrino can be generated but also diagonal terms with two dark fermions or two neutrinos. For a small mixing strength $\Theta$, the diagonal terms dominate over the absorption operators considered in this paper. Then one may expect much more $\gamma, mono$-jet, and mono-$Z$ events with the missing energy taken away by a pair of dark fermions $\chi\bar{\chi}$ or neutrinos ($\nu\bar{\nu}$). Consequently, the sensitivity can be even further improved with larger data sample. For very light dark fermion, $m_{\chi}\ll\sqrt{s}$, with mass $m_{\chi}$ being much smaller than the center-of-mass energy $\sqrt{s}$, the event rates for different combinations of final-state particles follow essentially the same spectrum shape and can simply scale up by a factor of $1/\Theta^{2}$. Since the event rates scale with the cut-off scale $\Lambda$ as $1/\Lambda^{4}$, a naive estimation of the sensitivity on $\Lambda$ would lead to a factor $1/\sqrt{\Theta}$ enhancement. However, we only focus on the absorption operators in the current study to be conservative and keep generality without too much model assumptions.
Furthermore, it is possible for the final-state dark fermion to have sizable decay width $\varGamma_{\chi}$. Especially, it may become visible at the collider detector if the dark fermion decays fast enough into the SM particles. It is then possible to distinguish the visible and invisible decay channels of the produced dark fermion at collider. However, adding visible decay channels would reduce the branching ratio of the invisible channels but would not reduce the total event rate of the $\gamma, mono$-jet, or mono-$Z$ events. Then the projected sensitivity can only be further enhanced rather than significantly reduced. For illustration, we consider the most conservative case with an invisible dark fermion in the current study.
However, the possibility of distinguishing the visible and invisible decay channels is an interesting extension of our current study. Although we would not get into too many details, it is useful to illustrate the probable parameter space. With absorption operators that involve quarks, the dark fermion can decay into a neutrino plus a pair of quarks, $\chi\to\nu+q+\bar{q}$. The total decay widths are given as,
$\varGamma=\left\{\frac{m_{\chi}^{5}}{1024\pi^{3}\varLambda_{S/P}^{4}},\frac{m_{\chi}^{5}}{256\pi^{3}\varLambda_{V/A}^{4}},\frac{3m_{\chi}^{5}}{128\pi^{3}\varLambda_{T}^{4}}\right\},$ (3.4)
for the five effective operator types. Notably, the total decay width has $m_{\chi}^{5}$ dependence. At a collider with a parton-level center-of-mass energy $\sqrt{\hat{s}}$, the typical decay length of the dark fermion is,
[…]
For a relatively light dark fermion, this simplifies to $L_{\chi}\approx\sqrt{\hat{s}}/2m_{\chi}\varGamma_{\chi}$. As a benchmark, we require the typical decay length to exceed one meter. Fig. 18 shows the regions where $L_{\chi}<1\,\text{m}$ for a typical LHC parton-level center-of-mass energy of $\sqrt{\hat{s}}=1\ \text{TeV}$.
Fig. 18: Typical decay length of the dark fermion in the $m_{\chi}$-$\varLambda$ plane. The parton level center of mass energy is chosen as $\sqrt{\hat{s}}=1\,{\rm TeV}$.
As one can see, the invisibility assumption is always valid for $m_{\chi}\lesssim 5\ \text{GeV}$ which covers the stable DM particle region with a tiny mass $m_{\chi}\lesssim 100$ MeV. Note that within this mass window, the direct decay channels are forbidden by the QCD color confinement. For a heavier dark fermion, the sensitivity should at least remain roughly the same if not better.